OpenStudy (anonymous):

A bag contains 6 purple marbles and 7 white marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability that the first marble is white and the second one is purple?

OpenStudy (anonymous):

choose the white marble first => 7 means(from 13 marbles) and then choose the purple one => 6 means (from 12 marbles 'cus the first one was drawn) all the sampling that can happen = 13*12 the propability = 7*6/13*12 = 42/156